OpenStudy (serena209):

Jada is trying to determine the number of t-shirts that were in inventory before she started her shift. During her shift, she sold eleven t-shirts, which left eighteen t-shirts in inventory. -Define a variable. -Write an equation Jada can use to determine the number of t-shirts that were in inventory before she started her shift. - Solve the equation showing your work.

OpenStudy (serena209):

its an essay

OpenStudy (serena209):

plzz help

OpenStudy (kobeni-chan):

b = number of shirts before her shift (what we're trying to find) She sold 11 shirts during her shift, and there were 18 shirts leftover, so you would subtract the 11 shirts from the original inventory (b) to get the leftover amount, 18. b - 11 = 18

OpenStudy (serena209):

thanks

OpenStudy (kobeni-chan):

No problem! You just have to solve the equation and show a step or two of how you solved it :)

OpenStudy (serena209):

ok can you help me with 2 more

OpenStudy (kobeni-chan):

Sure

OpenStudy (serena209):

If the graph shows the amount of money Mandy is aiming to have left in her wallet after a trip to the grocery store, could she have seven dollars left in her wallet? Explain your answer. A number line shows the numbers zero to ten. There is an open dot at six and an arrow extending from that open dot toward the left.

OpenStudy (serena209):

Match each equation with its solution. 1.21 - u = 21 2.y ^ 2 - 7 = 5 3. 4n - 8 = 40 4. g ^ 3= 1 =26 a.33 b.38 c.24 d.12

OpenStudy (serena209):

forget the other question

OpenStudy (kobeni-chan):

Sorry had to do something, I'm back now

OpenStudy (serena209):

its ok

OpenStudy (kobeni-chan):

Forget the other question, being the one with Mandy?

OpenStudy (serena209):

yes

OpenStudy (kobeni-chan):

Ok, it looks like a lot of the equations involve some subtracting. The first step would be to add these values to both sides of each equation. For example, in y^2 - 7 = 5, add 7 to both sides of the equal side to make y^2 = 12 Also, can you check if #1 is correct? Because u = 0 and that's not one of our choices :/

OpenStudy (kobeni-chan):

Can you simplify the equations or do you want me to help with that part?

OpenStudy (serena209):

what question are you answering

OpenStudy (serena209):

@kobeni-chan

OpenStudy (serena209):

your supposed to match them up

OpenStudy (kobeni-chan):

Yes, I'm walking you through how to solve it

OpenStudy (serena209):

ohh ok

OpenStudy (kobeni-chan):

Oh, did you double check that #1 was copied correctly?

OpenStudy (serena209):

yes

OpenStudy (kobeni-chan):

Well I know that 21 - 0 = 21, so let's leave that one for last. 2. y ^ 2 - 7 = 5 We simplified this to y^2 = 12 when we added 7 to both sides. The next step would be to find the square root of both sides, which gives us y = √12 Now the square root of 12 is about 3.46, which isn't one of the choices, but 12 is. So it might be that, but we should probably skip to the next one to see if we can get a definite answer for one of these equations.

OpenStudy (kobeni-chan):

Real quick, may I ask what subject of math this is for?

OpenStudy (serena209):

Math

OpenStudy (kobeni-chan):

Oh ok

OpenStudy (kobeni-chan):

I have a test of my own I'm working on, so tag me if you come back online and if you still need to finish the question

OpenStudy (serena209):

@kobeni-chan i just have to hurry